OpenStudy (anonymous):

Find the exact value of cotan(\sqrt{3}) I know how to compute exact values for the primary trig ratios, but how do you do so with their inverses?

OpenStudy (jamesj):

do you mean arctan(sqrt(3))?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I don't want the answer, rather I want to know what method we use to get it.

OpenStudy (jamesj):

Well remember sin(pi/3) = sqrt(3)/2 and cos(pi/3) = 1/2, hence ...

OpenStudy (anonymous):

What method we use to find it, and for any other inverse trig ratios, how would I get the exact value?

OpenStudy (jamesj):

In this case you should just learn the value of sin, cos and tan for all these angles x = 0, pi/6, pi/4, pi/3, pi/2

OpenStudy (jamesj):

and then you'll know this one.

OpenStudy (jamesj):

these and all the variations (pi/2 - x), (x + pi), etc. will then be available to you. Otherwise, frankly you're estimating. There are numerical methods to find arctan(a) for a given a.

OpenStudy (anonymous):

I don't get it. How would that help me get the inverse values?

OpenStudy (anonymous):

So for this, the answer is pi/6?

OpenStudy (jamesj):

Well, if you know cos(pi/4) = 1/sqrt(2) for instance, then you know that for the standard range of arccos, arccos(1/sqrt(2)) = pi/4

OpenStudy (jamesj):

Hence, for example, what is arctan(1)?

OpenStudy (jamesj):

sorry, you are calculating arctan(sqrt(3))

OpenStudy (jamesj):

then the answer is not pi/6.